A few pieces in the form of a file^{1}

I. Snoek

August 14, 1992

1paraphrasing E.Satie

I give here a few thingies I have made using that excellent program MED.Because I have a classical education(I am a musicology student) there are only classical pieces in it. I will give some extra information on all four of them:

1. ElGril.smp(El Grillo(=The Cricket)) This is a piece by the Dutch(!) composer Josquin des Prez(ca. 1440-1521).It is made in honour of a king.The underlying text is:

”El Grillo e buon cantore”(The Cricket is a good singer)

The Cricket in this case is off course the king. The musical form is a frottola.A characteristic of a frottola(pl. frottole) is the burden at the beginning and the end.A frottola is always secular.

The key or modality is mixolydian.

2. menuet.smp This is from the suite in B♭ major by J.S. Bach(1685-1750).There is not really much to tell about this piece.(I left the Trio-section out).

3. mmenuet.smp Another menuet bij Joseph Rheinsbürger(1674-1729)(That’s what it says in the sheet music) The key is G-minor.

4. VHH.smp This is the christmas song ”Von Himmel hoch” arranged for unaccompanied choir by J.S. Bach.It is supplied with the Magnificat.Note the long notes of the song above a very contrapuntal and mobile background.It is (I think) one of his masterpieces but unfortunately virtually unknown.

5. couperin2.smp

This is rondeau by Fr. Couperin,the famous 18th century harpsichord player.It is typically French because it is witty and light-footed.

Musique c’est pour faire plaisir.

This attitude towards music,by the way, has always been there in the French music up to the first halve of the twentieth century.Messiaen could be said the first not to have this attitude(His music and that of Boulez is totally not meant to listen to,it’s more of mathematical excercise).

The files with the extension *.smp are complete songs.(smp just means I’ve used samples). The file couperin2.smp needs OctaMEDPlayer, all other files only need MEDPlayer. If you have any suggestions or good(I mean good more than 1 octave) samples please mail to the address mentioned below or write to me personnally.(I don’t have an E-mail address)

E-mail:
cwquak@ruu.nl.cs
Normal mail:
Iede Snoek
Leuvenumseweg 70(from September 10th "v. Heemskercklaan 30")
3847 LD
Harderwijk
telephone:03410-19696